OpenStudy (anonymous):

What is the remainder of (x^3-6x-9x+3) divided by (x-3)

OpenStudy (anonymous):

(x^3-6x-9x+3) is it (x^3-6x^2-9x+3)?

OpenStudy (phi):

do you know how to divide ? or use synthetic division?

OpenStudy (anonymous):

its the first on you put @surdawi

OpenStudy (anonymous):

one

OpenStudy (anonymous):

and phi yes i do but idk how to divide like this

OpenStudy (anonymous):

ok thanks phi

OpenStudy (anonymous):

p(x)=(x^3-6x-9x+3 p(3) = (x^3-6x-9x+3

OpenStudy (anonymous):

Substitute x=3 in the eq and Find the remainder....Using..Remainder theorem..)

OpenStudy (anonymous):

27 - 18 - 27 + 3

OpenStudy (anonymous):

yahoo you see the (divided by) (x-3) right

OpenStudy (anonymous):

Can someone help

OpenStudy (anonymous):

A. -51 B. -51/x-3 C. -17/x D. -17/x-3

ganeshie8 (ganeshie8):

the question has a type, it must be : x^3-6x^2-9x+3 u do what yahoo said, u would get some answer thats there in options

OpenStudy (anonymous):

hold on and i will show you the ?

ganeshie8 (ganeshie8):

yea

OpenStudy (anonymous):

\[(x^{2}-6x-9x+3)\div(x-3)\]

ganeshie8 (ganeshie8):

if this question is correct, then we wont get any of the given answer choices

ganeshie8 (ganeshie8):

\((x^{2}-6x^2-9x+3)\div(x-3) \)

OpenStudy (anonymous):

agree with ganeshine

ganeshie8 (ganeshie8):

it should be like above, plz check wid ur teacher

OpenStudy (anonymous):

wait a sec.

OpenStudy (anonymous):

\[x^{3}-6x-9x+3)\div(x-3)\]

OpenStudy (anonymous):

now that question i just posted it correct

OpenStudy (anonymous):

so that means it is x^3 - 15x + 3

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

ok are the solutions A. -51 B. -51/x-3 or A. -15 B. -15/x-3

OpenStudy (anonymous):

A. -51 B. -51/x-3

OpenStudy (anonymous):

ok then it should be x^3-6*x^2-9*x+3 other wise it should be written in the book as x^3-15x+3 and the answer will be -51/x-3 if it is x^3-6x-9x+3 you will get -15/x-3

OpenStudy (anonymous):

ok thank you so much this one was really crazy

OpenStudy (anonymous):

it wasnt too crazy, you just forgot the ^2 :p

OpenStudy (anonymous):

lol i know :)